In January. Heart Orphanage received a sum of donation of $3000. It spent 12% sum received on books, 3/5 food and the rest on toiletries. in February, the amount of donation received increased. The orphanage spent the same amount of money on books but increased its spending on food by $420. It spent the remaining 25% on toiletries.
a) How much did the orphanage spend on toiletries in January?
b) What was the percentage increase in the amount of donation received?
Give your answer as a decimal correct to 1 decimal place

Answer by greenestamps(13200)



books in January: 12% of $3000 = $360
food in January: 3/5 of $3000 = $1800
toiletries in January: $3000-($360+$1800) = $840

books in February: $360 (same as January)
food in February: $1800+$420 = $2220
toiletries in February: 25% of the February total. So the $360+$2220 = $2580 February total spent on books and food was 75% of the February total. That means the amount spent on toiletries in February is 1/3 the total spent on books and food: 1/3 of $2580 = $860.

The total received in February was then $2580+$860 = $3440.

ANSWERS:
a) $840 on toiletries in January
b) % increase in donations from January to February: ((3440-3000)/3000)*100 = 14.7 to one decimal place